OpenStudy (watchmath):

Compute \[\lim_{x\to +\infty} x\left(\sqrt{x^2+1}-\sqrt[3]{x^3+1}\right)\]

OpenStudy (anonymous):

even with the "x" out front?

OpenStudy (watchmath):

yes, the problem is correct

OpenStudy (anonymous):

\[\lim_{x\to +\infty} x\times \frac{x}{x}\left(\sqrt{x^2+1}-\sqrt[3]{x^3+1}\right)\] \[ x^2 ( \sqrt{1 + \frac{1}{x^2}} + \sqrt[3]{1 + \frac{1}{x^3}})\] I think it must be \(\infty\)

OpenStudy (anonymous):

oh i meant with the x out front it seemed unlikely to be 0

OpenStudy (anonymous):

ishaan it is a minus sign. with a plus you would get infinity for sure

OpenStudy (anonymous):

it is minus

OpenStudy (turingtest):

lost a minus sign mertsj, right\[x^2 ( \sqrt{1 + \frac{1}{x^2}} - \sqrt[3]{1 + \frac{1}{x^3}})=\infty(0)\]so it is still indeterminate

OpenStudy (turingtest):

oh, ishaan, whoever it was lol

OpenStudy (anonymous):

Oh yeah, sorry. thanks satellite

OpenStudy (anonymous):

i will make a guess at a method, but it looks like a lot of work (algebra) i bet the usual gimmick will work of multiplying by the conjugate, but since you have powers of 1/2 and 1/3 you are going to have to use \[(a^6-b^6)=(a-b) (a+b) (a^2-a b+b^2) (a^2+a b+b^2)\]

OpenStudy (anonymous):

that is just a guess though

OpenStudy (anonymous):

*Bookmark*

OpenStudy (anonymous):

too much algebra for me, i can't do it

OpenStudy (mertsj):

You can. You just don't want to.

myininaya (myininaya):

i think i might have it

OpenStudy (watchmath):

hello myininaya. Still remember me? :)

myininaya (myininaya):

yes i do i'm getting the limit is infinity is that right?

myininaya (myininaya):

so i will show you what i did...

OpenStudy (anonymous):

@watchmath myininaya and i were just mentioning you the other day. sunday

OpenStudy (anonymous):

i said i missed your problems (not that i can do them) and no the answer is not infinity, but i cheated.

OpenStudy (watchmath):

no, the limit is finite

OpenStudy (anonymous):

i think the method i wrote would work if i could do the algebra

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x((x^2+1)^\frac{1}{2}-(x^3+1))^\frac{1}{3}}{1} \cdot \frac{(x^2+1)^\frac{1}{2}+(x^3+1)^\frac{1}{3}}{(x^2+1)^\frac{1}{2}+(x^3+1)^\frac{1}{3}}\] \[\lim_{x \rightarrow \infty}\frac{x(x^2+1-(x^3+1)^\frac{2}{3})}{(x^2+1)^\frac{1}{2}+(x^3+1)^\frac{1}{3}}\] \[\lim_{x \rightarrow \infty}\frac{x^3+x-x(x^3+1)^\frac{2}{3}}{(x^2+1)^\frac{1}{2}+(x^3+1)^\frac{1}{3}}\] So if I divide both top and bottom by x I get \[\lim_{x \rightarrow \infty}\frac{x^2+1-(x^3+1)^\frac{2}{3}}{(\frac{x^2+1}{x^2})^\frac{1}{2}+(\frac{x^3+1}{x^3})^\frac{1}{3}}\] \[\lim_{x \rightarrow \infty}\frac{x^2+1-(x^3+1)^\frac{2}{3}}{(1+\frac{1}{x^2})^\frac{1}{2}+(1+\frac{1}{x^3})^\frac{1}{3}}\] so we have that the bottom goes to 2 as x goes to positive infinity but the top goes to infinity so that is why i conclude the answer was infinity Maybe I should play with the top some more and hey how are you watchman lol i have missed your problems

OpenStudy (anonymous):

@myininaya your numerator is wrong

OpenStudy (anonymous):

middle terms don't add to zero because the powers are not the same. i think it the same gimmick will work, multiply something to clear the radicals, but since one power is 1/2 and the other is 1/3 i am pretty sure you have to use \[(a^6-b^6)=(a-b) (a+b) (a^2-a b+b^2) (a^2+a b+b^2)\]to clear them

myininaya (myininaya):

satellite i disagree i think my numerator is right

OpenStudy (anonymous):

oh i am sorry, yes it is right, but you still can't get the limit that way

OpenStudy (anonymous):

Here's a solution. Using the extension of the binomial theorem to arbitrary complex numbers, you have that \[(1+x^2)^\frac12=x+\frac12x^{-1}-\frac18x^{-3}+O(x^{-5})\] \[(1+x^3)^\frac13=x+\frac13x^{-2}+O(x^{-5})\] Thus, you have that \[\lim_{x\to \infty}x\left((1+x^2)^\frac12-(1+x^3)^\frac13\right)\] \[=\lim_{x\to\infty}x\left(\left(x+\frac12x^{-1}+O(x^{-3})\right)-\left(x+\frac13x^{-2}+O(x^{-5})\right)\right)\] \[=\lim_{x\to\infty}x\left(\frac12x^{-1}+O(x^{-2})\right)=\lim_{x\to\infty}\frac12+O(x^{-1})=\frac12\]

myininaya (myininaya):

imperialist is a showoff

OpenStudy (anonymous):

pretty damned nice

OpenStudy (watchmath):

imperialist is correct. Any elementary solution?

OpenStudy (anonymous):

I'm pretty sure satellite's method is also a good way to approach it, but as myininaya can attest, the algebra is not very friendly to say the least. Good job though myininaya!

OpenStudy (anonymous):

well i am still thinking of writing the difference of two sixth powers, the algebra is killing me

OpenStudy (anonymous):

i would say the above approach is much nicer, especially since i have filled up a piece of paper and probably have at least one mistake

OpenStudy (mertsj):

Why don't you brainiacs go help pre-algebra? He's been trying to get help most of the day.

myininaya (myininaya):

you know numerically my top approaches 1

myininaya (myininaya):

algebraically my bottom approaches 2

myininaya (myininaya):

if some how i can show without a numerical approach or graphical approach that the top approaches 1, then I would be dandy

OpenStudy (anonymous):

where is the pre-algebra question? i can't find it

OpenStudy (mertsj):

Updated 41 minutes ago.

myininaya (myininaya):

So how can I show that \[\lim_{x \rightarrow \infty}(x^2+1-(x^3+1)^\frac{2}{3})=1?\]

OpenStudy (anonymous):

Hmm x^2 + 1 - (x^3+1)^(2/3) can't we ignore 1?

myininaya (myininaya):

couldn't i say \[x^3+1 \approx x^3\]

myininaya (myininaya):

\[(x^3)^\frac{2}{3}=x^2\]

myininaya (myininaya):

x^2-x^2=0 we are left with 1

OpenStudy (anonymous):

Mhm, sounds good to me!

myininaya (myininaya):

i think it sounds awesome! :)

OpenStudy (watchmath):

If you can compute \[\lim_{x\to \infty} x\left(\sqrt[3]{x^3+1}-x\right)\] then you can compute the original problem :D.

OpenStudy (jamesj):

(For what it's worth, I think Imperialist has it. I can't see a proof that doesn't rely on some version of a power series expansion. I've learnt the hard way never to say never, so perhaps there something else more elementary out there. Nonetheless, sometimes elementary problems have complicated solutions.)

myininaya (myininaya):

James what do you think about what I was saying?

myininaya (myininaya):

We can say \[x^3 \approx x^3+1 \]

OpenStudy (jamesj):

The trouble is how exactly is that notation defined? Intuitively, what you've done is right. But can you be sure that it makes sense and doesn't lead to contradictions. The advantage of what Imperialist has done is all the notation he's used is very defined and can be in such a way that we can be sure of our conclusions.

OpenStudy (watchmath):

if you are expecting this problem to be answered elementary my suggestion is to split it into computing \(\lim x(\sqrt{x^2+1}-x)\) and \(\lim x(x-\sqrt{x^3+1})\)

OpenStudy (jamesj):

cube root in that second expression, right?

OpenStudy (watchmath):

yes, sorry for that

OpenStudy (jamesj):

that's a nice idea. wish I'd thought of it.